Analysis of test beam results (�le RP1 muons20.dat)

events total  10550

non-empty events 8220

events with at least 3 active planes per event 5359

events with at least 3 active planes per event after track selection 5224

Table 1: Events statistics.

The steps in the selection procedure

1) Reject all events with less than 3 active planes in any projection.

2) Apply track search and select hits (clusters) belonging to the recognized track.

3) Reject all events with less than 3 active planes in any projection.

Figure 1: The numbers of active planes per event. Left: u projection, right: v projection. Black histogram represents situation before step 1 while the red one before step 3.

Figure 2: Minimal number of active planes (the minimum is taken between u and v projection). Black histogram represents situation before step 1 while the red one before step 3.

Figure 3: 2D histogram of the numbers of active planes. Before step 1.

Figure 4: The difference between the numbers of active planes in u and v projection, in absolute value. Before step 1.

Figure 5: The histogram of the number of rejected hits (clusters) per event. During the step 2.

selected hits (clusters) 49410

rejected hits (clusters) 6124

Table 2: Hits statistics. The counts correspond to the entire run.